# Following the Human Thread in Social Navigation

### v1

@@ Abstract
We propose the first Social Dynamics Adaptation model (SDA) based on the robot’s state-action history to infer the social dynamics. We propose a two-stage Reinforcement Learning framework: the first learns to encode the human trajectories into social dynamics and learns a motion policy conditioned on this encoded information, the current status, and the previous action. Here, the trajectories are fully visible, i.e.,  assumed as privileged information. In the second stage, the trained policy operates without direct access to trajectories. Instead, the model infers the social dynamics solely from the history of previous actions and statuses in real-time. Tested on the novel Habitat 3.0 platform, SDA sets a novel state of the art (SoA) performance in finding and following humans.
The code is available at: https://github.com/L-Scofano/SDA.
Keywords:
Embodied AI Social Navigation Human Trajectories
**footnotetext: Authors contributed equally. This project was supported by PNRR MUR project PE0000013-FAIR. We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support.
Figure 1: We present our novel Social Dynamic Adaptation model (SDA). The framework involves two stages of training that allow the model to infer, given its past observations and actions, another agent’s Social Dynamics. In the first training stage, the model embeds the followed agent’s trajectory, which together with sensor perceptions compose the input to the model’s Social navigation Policy ($\pi$). The knowledge obtained from the human trajectory strongly helps the navigation policy in finding and following an agent, however, this information is often not available during deployment. In the second stage, SDA learns to adapt past statuses and actions, which are always available, to the first stage’s Social Dynamics embedding $\hat{z}$. As depicted in the figure, the status contains depth maps and BB detection of the person, if observable from the egocentric robot view. $\hat{z}$ is then paired with current observations as input to the frozen $\pi$.
@@ 1 Introduction
Traditional navigation techniques within Embodied Artificial Intelligence (EAI) have marked a crucial advancement by introducing robots into real-life environments. However, these techniques have primarily focused on agents traversing vacant spaces. Conversely, the significance of social navigation within EAI has steadily increased. Social navigation entails agents’ capacity to navigate human-centric environments while considering human movements and behaviours. These agents need to be able to locate, track, and interact with humans in a safe and socially acceptable manner. Previous studies predominantly characterized Social Navigation as a variation of PointGoal Navigation, wherein agents strive to reach specified destinations while considering human movements [48, 50, 32]. Habitat 3.0 [36], a significant breakthrough in EAI, introduces a lifelike environment seamlessly incorporating human avatars. This integration enables the investigation of human-agent interactions within a controlled, risk-free, dynamic environment. What sets Habitat 3.0 [36] apart is its ability to replicate complex scenarios where human intentions are constantly changing. Nevertheless, this dynamism also presents particular challenges, such as collision avoidance and achieving success in locating and following humans.
@@
Despite notable efforts in collision avoidance and safety, most existing methods for Social Navigation either rely on privileged information that is not available in real-world scenarios or do not adequately capture the social dynamics and norms of human behaviour. For instance, [48] and [50] use a GPS and compass sensor to provide the agent with perfect localization, which might be unrealistic even if using SLAM [31, 17] methods, whenever the human is not in the line of sight. While [32] and [52] do not account for the social factors that influence human behaviour. Therefore, this limitation hinders their practical applicability and adaptability to dynamic environments. [9], instead, models some social factors in the form of Proximity Tasks but fails to account for the cooperative nature that a social agent must possess, restricting itself to merely avoiding collisions with them. The current SoA model proposed in [36] necessitates privileged information, such as humanoid GPS, which offers polar coordinates, detailing the accurate distance and angle of the human from the robot, to attain high-performance outcomes. However, this requirement is highly impractical in real-world environments during inference.
@@
This paper proposes a novel Social Dynamics Adaptation model (SDA), shown in Fig. 2, that effectively solves the robot’s awareness of complex human behaviors, even temporarily losing sight of the person and fast robot motion. Specifically, the first stage trains a base policy considering human trajectories encoded into a latent vector. The latent vector is a low-dimensional, nonlinear projection of the human trajectories, and it is trained end-to-end with the base policy to extract the social factors that led the robots to choose better actions. The subsequent supervised adaptation stage regresses this latent vector using only the robot’s state and action history. Unlike previous methods, such as [48] and [52], which often rely on simulated privileged information or simplified social behavior models, our approach learns to respond to real-world dynamics in real time. In summary, SDA adapts and accounts for unpredictable human behaviors by exploiting privileged information during training and recovering this fundamental signal during deployment when it is often impractical to compute. Finally, the deployed robot features the motion policy, learned in the first stage, and the social dynamics, inferred from prior statuses and actions.
@@
Out of extensive benchmarking, SDA outperforms the approach proposed in Habitat 3.0 [36] and a second adapted best performer method [9] from Habitat 2.0 [43]. We conduct a thorough experimental evaluation on the core contribution of this work, i.e., learning to infer the social dynamics from (privileged) information about the person. Remarkably, out of ablative studies, we conclude that not only are human trajectories strong input information for the robot control policy, but they also make better supervision for inferring the social dynamics latent for the same policy. Other oracular information such as the humanoid GPS (direction and distance from the person-to-robot) are powerful sensors for the control policy but do not yield adaptable social dynamics to infer the human-robot-scene interaction.
@@ 2 Related Work
Recently, there has been a surge in exploring indoor navigation within an embodied framework [16]. This upswing has been facilitated primarily by the availability of large-scale datasets comprising 3D indoor environments [10, 42, 39] and simulators designed for simulating navigation within these dynamic 3D spaces [41, 42, 23]. Nevertheless, these simulators are not equipped to handle human entities within the environments, restricting the investigation to navigation tasks in scenarios where the agent functions independently or, at most, alongside humans simulated via static meshes that simulate movement [49, 52]. These simulated humans are treated as dynamic obstacles and lack compliance with any social construct. This constraint has been effectively addressed with Habitat 3.0 [36], the simulator used for this research. Habitat 3.0 introduces the capability to simulate the behaviours of humans engaging in tasks within dynamic environments, thus overcoming the limitations mentioned above. Furthermore, there is no longer a need to directly train social models in the real world, where there is a tangible risk of causing harm to actual humans. Thus, Habitat 3.0 establishes itself as a viable proxy for real-world adaptation of social models trained in simulation.
@@
Thanks to these simulators the realm of EAI has witnessed the introduction of numerous tasks [16], including PointGoal Navigation [48], ObjectGoal Navigation [4], Embodied Question Answering [47], and Vision and Language Navigation (VLN) [2, 24]. Various modular approaches [8, 12, 11, 38, 40] have been proposed to address the challenges of navigating through static, single-agent environments. These approaches utilize maps constructed from depth images and conduct path planning directly on these maps. However, these approaches are unsuitable in social settings where dynamic objects (humans) move within the environment. This is because humans are observable only within the agent’s field of view (FOV). Moreover, the agent must address the additional challenge of tracking and mapping. End-to-end RL-trained policies [46, 32, 51, 7, 50], should be adapted to learn also social clues similarly to [9], where the agent learned proxemics information about the humans moving in the environments thanks to two proximity tasks. In this paper, instead, we try to learn social behaviours directly by internally modeling the humanoid trajectories in the latent representation of the agent. Furthermore, differently from [9], the agent’s aim is no longer just avoiding collisions. Still, it involves locating this dynamically acting human and following them for a specified number of steps while maintaining a safe distance. This evolution of the task demands a heightened level of social comprehension from the agent, requiring the anticipation of the person’s intentions and the ability to trail them closely without compromising safety.
@@
Research in robotics, computer vision, and the analysis of human social behavior explored socially aware representations and models [30]. Extending from the realm of collision-free multi-agent navigation [6, 5, 15, 27] and navigation in dynamic environments [3], researchers have further expanded their investigations to encompass scenarios involving human presence [20, 18, 14, 29, 13].
The approach presented in [14] incorporates collision avoidance algorithms like CADRL [15] while introducing common-sense social rules. This integration aims to reduce uncertainty while minimizing the risk of encountering the Freezing Robot Problem [44]. Other works [18, 13] seek to model human-agent interaction by employing techniques such as spatiotemporal graphs [29].
Typically, these methods undergo testing in a minimal number of environments that offer complete knowledge about the human positions and velocities [13, 15], featuring simple obstacles and often assuming collaboration between moving agents[18, 13]. In contrast, our focus revolves around SocialNav within expansive indoor environments, characterized by partial knowledge about them, since the agent perceives the environment only through its sensors from an egocentric perspective, and no information about the velocity or position of the human is given to the agent. Our SDA addresses the missing information from the interacting human position, inferring it from the robot’s history of actions and status.
@@
Simulated environments [26, 45, 35] offer privileged information about the scene whose exploitation can be computationally intense or unfeasible during deployment. While navigating social environments, it is vital to take into consideration human behaviour[22]. Ideally, one would want to forecast people’s position for better path planning [33], but forecasting robot-person interactions is significantly slower [37] than navigation policies, hence being challenging to be considered for training or deployment. To overcome this problem, we leverage literature on system identification [1, 19] to infer the encoded privileged information during robot navigation. Once encoded in a latent space during the first training stage and used to train the primary policy, it is possible to asynchronously derive that same information from the state-action history [25, 28, 53], influenced by the signal we want to identify. Unlike previous works, we are the first to identify the social dynamics (under the form of human trajectories), with the intuition that modeling human behavior is fundamental for efficient human-robot collaboration.
@@ 3 Methodology
This section introduces a novel framework, SDA , for social navigation that incorporates human trajectories into the sequential decision-making process. The first stage of our approach focuses on encoding human trajectories into a latent social context to represent the social dynamics that are functional to the agent motion policy. In the second stage, we introduce the Adapter module, which enables the estimation of social information from the agent’s past behavior. Recovering this signal allows the robot to operate in the real world without explicitly representing human behavior. We elucidate the training process, trajectory modeling, and optimization techniques utilized in our approach, highlighting its effectiveness in addressing the challenges of social navigation.
@@
The problem involves the task of locating and following a humanoid in motion within a household environment, maintaining a distance of 1 to 2 meters for at least $k$ steps [36].
With the status $x_{t}$, we represent the agent’s “perception” at time $t$. We exploit depth images from different cameras placed on the robot and a preprocessed version containing a humanoid detection bounding box. These perceptions are processed with a ResNet [21] before being fed to the recurrent policy network, selecting the best action $a_{t}$ at time $t$. We use Decentralized Distributed Proximal Policy Optimization (DD-PPO) [48] to iteratively improve the agent’s policy while maximizing rewards derived from interactions with several environments executed in parallel, ensuring stability through controlled policy updates and lower training times.
The pipeline described above can be considered a baseline implementation that does not contain privileged information but relies only on the robot’s onboard sensors. To address the social aspect required by the task, we consider additional details on the humanoid, e.g., “social dynamics”, defined as $e_{t-N:t-1}$. In the following sections, we describe how exploiting this information at training time can improve performance during deployment when it is absent.
3.1 Stage 1: Social Policy
@@
where $r_{t}(\theta)$ is the ratio of the probability of action $a_{t}$ under the current policy and the previous one that is being executed for gathering data. $\hat{A}_{t}$ represents the advantage function at time $t$, guiding the policy towards actions that yield higher expected rewards.
Defining what information is considered “privileged” and why is essential. In our context, it refers to detailed knowledge about the humanoid in the environment, such as the exact position of humanoids defining a trajectory (traj.), of the relative position with respect to the agent often denoted as humanoid GPS [36](hGPS). We can easily gather this information in simulated environments. However, collecting them in real-life scenarios is often impractical. This distinction is crucial as it highlights the challenge of transferring learned policies from simulation to the real world.
3.2 Stage 2: Social Dynamics Regression
We aim to extract and exploit social cues directly from the robot’s perception and eliminate the need for auxiliary devices like GPS trackers on humanoids. Inspired by [25], we introduce the “social dynamics” module (Adapter), parametrized by an MLP $\psi$ that takes as input the recent history of the robot’s states $x_{t-N:t-1}$ and actions $a_{t-N:t-1}$ to generate a new latent vector $\hat{z}_{t}$:
@@ 3.2 Stage 2: Social Dynamics Regression
During this process, we optimize the Adapter, MLP, with a supervised regression objective, Mean Squared Error (MSE), to recover the original information contained in $z_{t}$ that we compute relying on the preferential information trajectory:
$\text{MSE}(\hat{z}_{t},z_{t})=\|\hat{z}_{t}-z_{t}\|_{2}^{2}$ (6)
Once we finalize the Adapter training, instead of relying on the privileged information, we can depend upon the robot’s states $x_{t-N:t-1}$ and actions $a_{t-N:t-1}$ to generate $\hat{z}_{t}$, serving as an estimate of the actual latent social dynamics vector $z_{t}$. Doing so enables the agent to estimate social dynamics online, improving its performance in dynamic environments and enhancing its social navigation capabilities, freeing it from dependence upon external sensors.
3.3 Training details
In our training process, we encode social dynamics using a ResNet, which we train from scratch. The trajectory encoder $\mu$ consists of a 3-layer MLP, and the output $z_{t}$ has a dimensionality of 128. The Adapter $\psi$ comprises alternating spatial and temporal MLP layers, with the output $\hat{z}_{t}$ matching the dimensionality of $z_{t}$. We jointly train the policy $\pi$, the perceptions encoder (ResNet), and the social dynamics (Trajectory Encoder) encoder $\mu$ during the first stage. In the second stage, everything is frozen except the Adapter.
In Stage 1, we utilize DD-PPO [48] for 250 million steps across 24 environments, following the training protocol presented in [36]. The training process takes approximately four days. In Stage 2, we employ supervised learning for 5 million steps across the same environments. The learning process lasts around 2 hours. Both stages utilize 4 A100 GPUs for efficient computation.

### v2

@@ Abstract
We present the first Social Dynamics Adaptation model (SDA) based on the robot’s state-action history to infer the social dynamics. We propose a two-stage Reinforcement Learning framework: the first learns to encode the human trajectories into social dynamics and learns a motion policy conditioned on this encoded information, the current status, and the previous action. Here, the trajectories are fully visible, i.e.,  assumed as privileged information. In the second stage, the trained policy operates without direct access to trajectories. Instead, the model infers the social dynamics solely from the history of previous actions and statuses in real-time. Tested on the novel Habitat 3.0 platform, SDA sets a novel state-of-the-art (SotA) performance in finding and following humans.
The code can be found at https://github.com/L-Scofano/SDA.
Figure 1: We present our novel Social Dynamic Adaptation model (SDA). The framework involves two stages of training that allow the model to infer, given its past observations and actions, another agent’s Social Dynamics. In the first training stage, the model embeds the followed agent’s trajectory, which, together with sensor perceptions, compose the input to the model’s Social navigation Policy ($\pi$). The knowledge obtained from the human trajectory strongly helps the navigation policy in finding and following an agent. However, this information is often not available during deployment. In the second stage, SDA learns to adapt past statuses and actions, which are always available, to the first stage’s Social Dynamics embedding $\hat{z}$. As depicted in the figure, the status contains depth maps and BB detection of the person, if observable from the egocentric robot view. $\hat{z}$ is then paired with current observations as input to the frozen $\pi$.
@@ 1 Introduction
Traditional navigation techniques within Embodied Artificial Intelligence (EAI) have marked a crucial advancement by introducing robots into real-life environments. However, these techniques have primarily focused on agents traversing vacant spaces. Conversely, the significance of social navigation within EAI has steadily increased. Social navigation entails agents’ capacity to navigate human-centric environments while considering human movements and behaviours. These agents need to be able to locate, track, and interact with humans in a safe and socially acceptable manner. Previous studies predominantly characterized Social Navigation as a variation of PointGoal Navigation, wherein agents strive to reach specified destinations while considering human movements (Wijmans et al., 2019b; Ye et al., 2021; Partsey et al., 2022). Habitat 3.0 (Puig et al., 2024), a significant breakthrough in EAI, introduces a lifelike environment seamlessly incorporating human avatars. This integration enables investigating human-agent interactions within a controlled, risk-free, dynamic environment. What sets Habitat 3.0 (Puig et al., 2024) apart is its ability to replicate complex scenarios where human intentions are constantly changing. Nevertheless, this dynamism also presents particular challenges, such as collision avoidance and achieving success in locating and following humans. Finding a person and following them is relevant to human-robot collaboration. For instance, in search-and-rescue operations, a robot may need to find and reliably follow a human responder through unpredictable and hazardous terrain to assist, carry equipment, or relay critical information. Similarly, in assistive robotics, such as elder care, robots must follow caregivers or patients across different rooms in a home, adapting seamlessly to changes in speed, direction, and environment. The complexity of the task stems from difficult long-term human motion prediction and the possible changes of their intentions, from having to navigate unknown and dynamic environments, and from strict safety requirements.
@@
Despite notable efforts in collision avoidance and safety, most existing methods for Social Navigation either rely on privileged information that is not available in real-world scenarios or do not adequately capture the social dynamics and norms of human behaviour. For instance, (Wijmans et al., 2019b) and (Ye et al., 2021) use a GPS and compass sensor to provide the agent with perfect localization, which might be unrealistic even if using SLAM (Mur-Artal et al., 2015; Engel et al., 2014) methods, whenever the human is not in the line of sight. While (Partsey et al., 2022) and (Yokoyama et al., 2022) do not account for the social factors that influence human behaviour. Therefore, this limitation hinders their practical applicability and adaptability to dynamic environments. (Cancelli et al., 2023), instead, models some social factors in the form of Proximity Tasks but fails to account for the cooperative nature that a social agent must possess, restricting itself to merely avoiding collisions with them. The current SotA model proposed in (Puig et al., 2024) necessitates privileged information, such as humanoid GPS, which offers polar coordinates, detailing the accurate distance and angle of the human from the robot, to attain high-performance outcomes. However, this requirement is highly impractical in real-world environments during inference.
@@
This paper proposes a novel Social Dynamics Adaptation model (SDA), shown in Fig. 2, that effectively solves the robot’s awareness of complex human behaviors, even temporarily losing sight of the person and fast robot motion. Specifically, the first stage trains a base policy considering human trajectories encoded into a latent vector. The latent vector is a low-dimensional, nonlinear projection of the human trajectories, and it is trained end-to-end with the base policy to extract the social factors that led the robots to choose better actions. The subsequent supervised adaptation stage regresses this latent vector using only the robot’s state and action history. Unlike previous methods, such as (Wijmans et al., 2019b) and (Yokoyama et al., 2022), which often depend on simulated privileged information or simplified social behavior models, our approach adapts to dynamics resembling real-world conditions in real-time. In summary, SDA adapts and accounts for unpredictable human behaviors by exploiting privileged information during training and recovering this fundamental signal during deployment when it is often impractical to compute. Finally, the deployed robot features the motion policy, learned in the first stage, and the social dynamics, inferred from prior statuses and actions.
@@
Out of extensive benchmarking, SDA outperforms the approach proposed in Habitat 3.0 (Puig et al., 2024) and a second adapted best-performing method (Cancelli et al., 2023) from Habitat 2.0 (Szot et al., 2021). We conduct a thorough experimental evaluation of the core contribution of this work—learning to infer social dynamics from (privileged) information about the person. Although our primary focus is on algorithm development, we also seek to improve the robustness of our method for potential real-world applications. To this end, we conduct experiments that bridge the gap to real-world scenarios by introducing noise to the input, reducing the refresh rate of the sensors, and modifying the simulator to reflect more realistic human behavior. Our ablative studies reveal that human trajectories are not only strong input information for the robot control policy but also provide better supervision for inferring the social dynamics latent to the same policy. Other oracular information, such as the humanoid GPS (direction and distance from the person to the robot), serves as powerful sensors for the control policy but does not facilitate adaptable social dynamics for inferring human-robot-scene interactions.
@@ 2 Related Work
Recently, there has been a surge in exploring indoor navigation within an embodied framework (Deitke et al., 2022). This upswing has been facilitated primarily by the availability of large-scale datasets comprising 3D indoor environments (Chang et al., 2017; Shen et al., 2021; Ramakrishnan et al., 2021) and simulators designed for simulating navigation within these dynamic 3D spaces (Savva et al., 2019; Shen et al., 2021; Kolve et al., 2017). Nevertheless, these simulators are not equipped to handle human entities within the environments, restricting the investigation to navigation tasks in scenarios where the agent functions independently or, at most, alongside humans simulated via static meshes that simulate movement (Xia et al., 2020; Yokoyama et al., 2022). These simulated humans are treated as dynamic obstacles and lack compliance with any social construct. This constraint has been effectively addressed with Habitat 3.0 (Puig et al., 2024), the simulator used for this research. Habitat 3.0 introduces the capability to simulate the behaviours of humans engaging in tasks within dynamic environments, thus overcoming the limitations mentioned above.
@@
Thanks to these simulators the realm of EAI has witnessed the introduction of numerous tasks (Deitke et al., 2022), including PointGoal Navigation (Wijmans et al., 2019b), ObjectGoal Navigation (Batra et al., 2020), Embodied Question Answering (Wijmans et al., 2019a), and Vision and Language Navigation (VLN) (Anderson et al., 2018; Krantz et al., 2020). Various modular approaches (Campari et al., 2022; Chaplot et al., 2020b; a; Ramakrishnan et al., 2022; Raychaudhuri et al., 2024) have been proposed to address the challenges of navigating through static, single-agent environments. These approaches utilize maps constructed from depth images and conduct path planning directly on these maps. However, these approaches are unsuitable in social settings where dynamic objects (humans) move within the environment. This is because humans are observable only within the agent’s field of view (FOV). Moreover, the agent must address the additional challenge of tracking and mapping. End-to-end RL-trained policies (Wani et al., 2020; Partsey et al., 2022; Ye et al., 2020; Campari et al., 2020; Ye et al., 2021), should be adapted to learn also social clues similarly to (Cancelli et al., 2023), where the agent learned proxemics information about the humans moving in the environments thanks to two proximity tasks. In this paper, instead, we try to learn social behaviours directly by internally modeling the humanoid trajectories in the latent representation of the agent. Furthermore, differently from (Cancelli et al., 2023), the agent’s aim is no longer just avoiding collisions. Still, it involves locating this dynamically acting human and following them for a specified number of steps while maintaining a safe distance. This evolution of the task demands a heightened level of social comprehension from the agent, requiring the anticipation of the person’s intentions and the ability to trail them closely without compromising safety.
@@
Research in robotics, computer vision, and the analysis of human social behavior explored socially aware representations and models (Möller et al., 2021). Extending from the realm of collision-free multi-agent navigation (Berg et al., 2011; Van den Berg et al., 2008; Chen et al., 2017b; Long et al., 2017) and navigation in dynamic environments (Aoude et al., 2013), researchers have further expanded their investigations to encompass scenarios involving human presence (Guzzi et al., 2013; Ferrer et al., 2013; Chen et al., 2017a; Lu et al., 2022; Chen et al., 2019).
The approach presented in (Chen et al., 2017a) incorporates collision avoidance algorithms like CADRL (Chen et al., 2017b) while introducing common-sense social rules. This integration aims to reduce uncertainty while minimizing the risk of encountering the Freezing Robot Problem (Trautman & Krause, 2010). Other works (Ferrer et al., 2013; Chen et al., 2019) seek to model human-agent interaction by employing techniques such as spatiotemporal graphs (Lu et al., 2022). Typically, these methods undergo testing in a minimal number of environments that offer complete knowledge about the human positions and velocities (Chen et al., 2019; 2017b), featuring simple obstacles and often assuming collaboration between moving agents(Ferrer et al., 2013; Chen et al., 2019). In contrast, our focus revolves around SocialNav within expansive indoor environments, characterized by partial knowledge about them, since the agent perceives the environment only through its sensors from an egocentric perspective, and no information about the velocity or position of the human is given to the agent. Our SDA addresses the missing information from the interacting human position, inferring it from the robot’s history of actions and status.
@@
Simulated environments (Li et al., 2021; Tsoi et al., 2022; Puig et al., 2018) offer privileged information about the scene whose exploitation can be computationally intense or unfeasible during deployment. While navigating social environments, it is vital to take into consideration human behaviour (Kivrak et al., 2021). Ideally, one would want to forecast people’s position for better path planning (Patle et al., 2019), but forecasting robot-person interactions is significantly slower (Rahman et al., 2023) than navigation policies, hence being challenging to be considered for training or deployment. To overcome this problem, we leverage literature on system identification (Ahmed & Qin, 2009; Guo et al., 2016) to infer the encoded privileged information during robot navigation. Once encoded in a latent space during the first training stage and used to train the primary policy, it is possible to asynchronously derive that same information from the state-action history (Kumar et al., 2021; Loquercio et al., 2023; Zhang et al., 2023; Liang et al., 2024; Kumar et al., 2022; Qi et al., 2023), influenced by the signal we want to identify. Unlike previous works, we are the first to identify the social dynamics (under the form of human trajectories), with the intuition that modeling human behavior is fundamental for efficient human-robot collaboration.
@@ 3 Methodology
This section introduces SDA : a novel framework for social navigation that incorporates human trajectories into the sequential decision-making process. The first stage of our approach focuses on encoding human trajectories into a latent social context to represent the social dynamics that are functional to the agent motion policy. In the second stage, we introduce the Adapter module, which enables the estimation of social information from the agent’s past behavior. Recovering this signal allows the robot to operate without explicitly representing human behavior. We detail the training process, trajectory modeling, and optimization techniques utilized in our approach, highlighting its effectiveness in addressing the challenges of social navigation.
@@
The problem requires to locate and follow a humanoid in motion within an indoor environment, maintaining a distance of 1 to 2 meters for at least $k$ steps (Puig et al., 2024). With the status $x_{t}$, we represent the agent’s “perception” at time $t$. We exploit depth images from different cameras placed on the robot and a preprocessed version containing a humanoid detection bounding box. These perceptions are processed with a ResNet (He et al., 2016) before being fed to the recurrent policy network, selecting the best action $a_{t}$ at time $t$. We use Decentralized Distributed Proximal Policy Optimization (DD-PPO) (Wijmans et al., 2019b) to iteratively improve the agent’s policy while maximizing rewards derived from interactions with several environments executed in parallel, ensuring stability through controlled policy updates and lower training times. The pipeline described above can be considered a baseline implementation that does not contain privileged information but relies only on the robot’s onboard sensors. To address the social aspect required by the task, we consider additional details on the humanoid, e.g., “social dynamics”, defined as $e_{t-N:t-1}$, where $N$ refers to the trajectory length. $e_{t-n}$ represents the position $n$ steps before time $t$, and $e_{t-N}$ refers to the earliest position in the trajectory under consideration. The notation $e_{t-N:t-1}$ is used as a shorthand to indicate the complete trajectory from $N$ steps in the past to the current time, representing absolute x-y coordinates. In the following sections, we describe how exploiting this information at training time can improve performance during deployment when it is absent.
Stage 1: Social Policy.
@@
where $r_{t}(\theta)$ is the ratio of the probability of action $a_{t}$ under the current policy and the previous one that is being executed for gathering data. $\hat{A}_{t}$ represents the advantage function at time $t$, guiding the policy towards actions that yield higher expected rewards. Defining what information is considered “privileged” and why is essential. In our context, it refers to detailed knowledge about the humanoid in the environment, such as the exact position of humanoids defining a trajectory (traj.), or the relative position with respect to the agent often denoted as humanoid GPS (Puig et al., 2024)(hGPS). We can easily gather this information in simulated environments. However, collecting them in real-life scenarios is often impractical. This distinction is crucial as it highlights the challenge of transferring learned policies from simulation to the real world.
Stage 2: Social Dynamics Regression.
We aim to extract and exploit social cues directly from the robot’s perception and eliminate the need for auxiliary devices like GPS trackers on humanoids. Inspired by (Kumar et al., 2021), we introduce the “social dynamics” module (Adapter), parametrized by an MLP $\psi$ that takes as input the recent history of the robot’s states $x_{t-N:t-1}$ and actions $a_{t-N:t-1}$ to generate a new latent vector $\hat{z}_{t}$:
@@
During this process, we optimize the Adapter, MLP, with a supervised regression objective, Mean Squared Error (MSE), to recover the original information contained in $z_{t}$ that we compute relying on the preferential information trajectory, $\text{MSE}(\hat{z}_{t},z_{t})=\|\hat{z}_{t}-z_{t}\|_{2}^{2}$. Once we finalize the Adapter training, instead of relying on the privileged information, we can depend upon the robot’s states $x_{t-N:t-1}$ and actions $a_{t-N:t-1}$ to generate $\hat{z}_{t}$, serving as an estimate of the actual latent social dynamics vector $z_{t}$. Doing so enables the agent to estimate social dynamics online, improving its performance in dynamic environments and enhancing its social navigation capabilities, freeing it from dependence upon external sensors.